A computer chip manufacturer spent $2,540,000 to develop a special-purpose molding machine. The machine has been used for one year and is expected to be obsolete after an additional 3 years. The company uses straight-line (SLN) depreciation for this machine.

At the beginning of the second year, a machine salesperson offers a new, vastly more efficient machine. This machine will cost $2,040,000, reduce annual cash manufacturing costs from $1,840,000 to $1,040,000, and have zero disposal value at the end of 3 years. Management has decided to use the double-declining-balance (DDB) depreciation method for tax purposes for this machine if purchased. (Note: Make sure to switch to SLN depreciation in year 3 to ensure that the entire cost of the asset is written off. You may find it useful to use the VDB function in Excel to calculate depreciation charges.)

The old machine’s salvage value is $304,000 now and is expected to be $54,000 three years from now; however, no salvage value is provided in calculating straight-line (SLN) depreciation on the old machine for tax purposes. The firm’s income tax rate is 45%. The firm desires to earn a minimum after-tax rate of return of 8%. (Use Table 1 and Table 2.) (Do not round intermediate calculations.)

Required:
Note: Use the PV and NPV functions in Excel to calculate all present value amounts.
1. What is the present value of tax savings associated with depreciating the existing machine (using the straight-line method)? (Round your final answer to the nearest whole dollar.)
2. What is the present value of tax savings associated with depreciating the new machine using the double-declining-balance method? Use the VDB built-in function in Excel to calculate depreciation deductions. (Round your final answer to the nearest whole dollar.)
3. What is the present value of net after-tax cost associated with the existing machine? (Hint: there will be three items to consider.) (Round your final answer to nearest whole dollar amount.)
4. What is the present value of the net after-tax cost of using the replacement (new) machine? (Round your final answer to the nearest whole dollar.)
5. What is the estimated net present value (NPV) of the decision to replace the existing machine with the new machine. (Round your final answer to the nearest whole dollar.)

1. Present value not attempted
2. Present value not attempted
3. Present value not attempted
4. Present value not attempted
5. Estimated net present value
Answer #1

FIRST 4 QUESTIONS ANSWERED

Ans-1

Present value of remaining Tax benefits on the existing machine :

Depreciation each year under SLN Method = 2540000/4 = 635000

The PV of the Tax benefits still remaing is as hereunder (using 8% rate)

Year Depreciation Tax Savings PV of Tax Savings (A) (B) = (A)*45% (C) = (B)/(1+8%)^n 2 635000 285750 264583 3 635000 285750 2

n = year lapse on which benefit will flow.Hence for Year 2 it will be 1 , Year 3 it will be 2 and Year 4 it will be 3

Salvage value not considered

Ans-2

PV of tax benefit of depreciation of new machine

2040000 New Machine Inititial Cost Salvage Value Life in Years 0 3 Year-2 Year-3 TOTAL VDB Depreciation Tax Benefit PV of Tax

Formula form used in above table

с D E F G 2040000 0 3 9 New Machine 10 Inititial Cost 11 Salvage Value 12 Life in Years 13 14 VDB Depreciation 15 Tax Benefit

Ans-3

NET COST ON EXISTING MACHINE Book Value Less:PV of Tax benefits on depreciation Less:PV of savings in manufacturing costs Net

Ans-4

There will be a net positive flow of 1148212 on new Machine as show in table below:

NET COST ON NEW MACHINE Purchase Cost of new Machine Less:Proceeds from old machine Less:PV of Tax benefits on depreciation L

Working-1

Year Savings PV of Savings 1 800000 740741 800000 685871 WN 800000 635066 2061678